Order: Nitroglycerin 50mg in 150 ml D5W over 24 hours by the infusion pump. What do you set your rate at? Order: R...

  1. Order: Nitroglycerin 50mg in 150 ml D5W over 24 hours by the infusion pump. What do you set your rate at?

  1. Order: Regular Insulin 20 units/hr

Supply: 250 units/250ml

What will your rate be (ml/hr)?

  1. The total volume to be infused is 1000ml. The rate is 75 ml/hr. How long will it take the infusion to run?
Answer #1

Ans) Order: Nitroglycerin 50mg in 150 ml D5W over 24 hours by the infusion pump.

ml/hr = Total volume to be infused ÷ time in hours

= 150/24 = 6.25 ml/hr.

Order: Regular Insulin 20 units/hr

Supply: 250 units/250ml

What will your rate be (ml/hr)?

The total volume to be infused is 1000ml. The rate is 75 ml/hr.

ml/hr = Total volume to be infused ÷ time in hrs

75 = 1000 ÷ time

Time = 1000/75 = 13.33 hrs ~ 13 hrs.